A pizza shop owner wishes to find the 95% confidence interval of the true mean cost of a large plain pizza. How large should the sample be if she wishes to be accurate to within $0.13? A previous study showed that the standard deviation of the price was $0.4.
Answer by Boreal(15235)

the margin of error should be +/-0.13, so 0.13=z*s/sqrt(n)
1.96*s/ sqrt (n)=0.13
1.96*0.4=0.13* sqrt(n), after multiplying through
divide through by 0.13 and square both sides
n=36.37 which rounds up to 37. That is unlikely to be changed by using a t-value